# O. Ragniscoa, F.Zullob

### v1

@@ 4.2 The case of an even Kernel
$\epsilon(x)=\frac{1}{\sqrt{2L}}\tilde{\epsilon}(x),\quad\eta(x)=-\eta(-x)$ (4.24)
@@
$\dot{Q}=\int_{-\infty}^{\infty}dx\epsilon(x)\exp[P\epsilon(x)+Q\eta(x)+\kappa(x)]$ (4.25)
@@
$S(P,x,L)=\int_{-L}^{L}\epsilon(\xi)\exp(P\epsilon(\xi))f(xs(\xi))d\xi,$ (4.26)
where $f(x)$ is a continuous function of $x$ and $\epsilon(\xi)$ is given in (4.24). Then, for any finite values of $x$ one has
$\lim_{L\to\infty}S(P,x,L)=Pf(0).$ (4.27)
Proof: by taking the Taylor expansion of the exponential we get
$S(P,x,L)=\int_{-L}^{L}\epsilon(\xi)(1+P\epsilon(\xi)+\frac{P^{2}}{2}\epsilon(\xi)+\ldots)f(xs(\xi))d\xi.$ (4.28)
@@
$F(x,L)=\frac{1}{2L}\int_{-L}^{L}\left(f(xs(\xi))-f(0)\right)d\xi$ (4.30)
@@
$\lim_{L\to\infty}\frac{1}{2L}\int_{-L}^{L}\left(f(xs(\xi))-f(0)\right)d\xi=\lim_{L\to\infty}\frac{f(xs(L))+f(xs(-L))-2f(0)}{2}=0,$ (4.32)
where in the last equation we use the fact that $s(\xi)\in S^{\infty}$.. Put in another way, in the Taylor series of the integral of $f(xs(\xi))$ (4.29) around $x=0$, only the first term $f(0)$ gives a finite contribution, all the others approaches zero in the limit $L\to\infty$.
From the previous result it follows that all the powers of $\epsilon$ greater than two doesn’t give a contribution in the limit $L\to\infty$.
Let us look at the term linear in $\epsilon$ in (4.26). One has
$\int_{-L}^{L}\epsilon(\xi)f(xs(\xi))d\xi=\frac{1}{\sqrt{2L}}\int_{0}^{L}\tilde{\epsilon}\left(f(xs(\xi))-f(xs(-\xi))\right)d\xi$ (4.33)
@@
$\frac{1}{\sqrt{2L}}\int_{0}^{L}\tilde{\epsilon}\Big(f(xs(\xi))-f(0)\Big)d\xi-\int_{0}^{L}\tilde{\epsilon}\Big(f(xs(-\xi))-f(0)\Big)d\xi$ (4.34)
@@
$\begin{split}&|\int_{-L}^{L}\epsilon(\xi)f(xs(\xi))d\xi|\leq\frac{C|x|}{\sqrt{2L}}\int_{0}^{L}\left(|s(\xi)-s(-\xi)|\right)d\xi\end{split}$ (4.35)
@@
Corollary 4.2
The equation of motion (4.25) reduces, in the limit $L\to\infty$, to
$\dot{Q}=P.$ (4.36)
@@
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(P\epsilon+Q\eta+\kappa)d\xi.$ (4.37)
@@
$T(P,x,L)=\int_{-L}^{L}\eta(\xi)\exp(P\epsilon(\xi))f(xs(\xi))d\xi,$ (4.38)
where $f(x)$ is a continuous function of $x$ and $\epsilon(\xi)$ is given in (4.24). Then, for any finite values of $x$ one has
$\lim_{L\to\infty}T(P,x,L)=\int_{-\infty}^{\infty}\eta(\xi)f(xs(\xi))d\xi$ (4.39)
Proof: For $\xi>\xi_{0}$, since $s(\xi)\to 0$ for $\xi\to\infty$, by continuity one has $|f(xs(\xi))-f(0)|<\delta(x)$ or $|f(xs(\xi))|<|f(0)|+\delta(x)$ So one has
$\begin{split}&|\int_{-\infty}^{\infty}\eta(\xi)f(xs(\xi))d\xi|=|\int_{|t|<t_{0}}\eta(\xi)f(xs(\xi))d\xi+\int_{|t|\geq t_{0}}\eta(\xi)f(xs(\xi))d\xi|\leq\\ &\leq\int_{|t|<t_{0}}|\eta(\xi)f(xs(\xi))|d\xi+(|f(0)|+\delta(x))\int_{|t|\geq t_{0}}|\eta(\xi)|d\xi,\end{split}$ (4.40)
i.e. the integral $\int_{-\infty}^{\infty}\eta(\xi)f(xs(\xi))d\xi$ is finite since $\eta(\xi)$ is in $S^{\infty}$. It follows that if one considers the Taylor expansion of $\exp(P\epsilon)$ in the integral (4.38), only the term proportional to $\epsilon^{0}$ survives in the limit $L\to\infty$ and we get (4.39). $\square$
@@
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi.$ (4.41)
@@
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(\kappa(\xi))\sinh(Q\eta(\xi))d\xi.$ (4.42)
@@
$\mathcal{H}=\frac{P^{2}}{2}-Q+\int_{-\infty}^{\infty}\exp(\kappa(\xi))\left(\cosh(Q\eta(\xi))-1\right)d\xi$ (4.43)
@@ 4.3 The case of an odd Kernel
$\epsilon(x)=\frac{1}{\sqrt{2L}}\tilde{\epsilon}(x),\quad\eta(x)=\eta(-x)$ (4.44)
@@
$\lim_{L\to\infty}\int_{-L}^{L}\epsilon(\xi)d\xi=o(\sqrt{L}).$ (4.45)
@@
$M(P,x,L)=\int_{-L}^{L}\epsilon(\xi)\exp(P\epsilon(\xi))f(xs(\xi))d\xi,$ (4.46)
where $f(x)$ is a continuous function of $x$ and $\epsilon(\xi)$ is given in (4.44) and satisfies (4.45). Then, for any finite values of $x$ one has
$\lim_{L\to\infty}M(P,x,L)=Pf(0).$ (4.47)
Proof: we take the Taylor series of the exponential in (4.46) one has
$M(P,x,L)=\int_{-L}^{L}\epsilon(\xi)\left(1+P\epsilon(\xi)+\frac{P^{2}}{2}\epsilon(\xi)^{2}+\ldots\right)f(xs(\xi))d\xi.$ (4.48)
@@
$P\int_{-L}^{L}\epsilon(\xi)^{2}f(xs(\xi))d\xi=\frac{1}{2L}\int_{-L}^{L}f(xs(\xi))d\xi.$ (4.49)
Exactly as for the equation (4.29), the previous expressions equals $Pf(0)$. It remains only the term linear in $\epsilon$ to look at, since all the powers of $\epsilon$ greater than $2$ in the Taylor expansion (4.48) do not contribute in the limit $L\to\infty$. The linear term in $\epsilon$ is
$\int_{-L}^{L}\epsilon(\xi)f(xs(\xi))d\xi=\frac{1}{\sqrt{2L}}\int_{-L}^{L}\tilde{\epsilon}(\xi)f(xs(\xi))d\xi.$ (4.50)
Due to the condition (4.45), since $f(xs(\xi))\to f(0)$ as $\xi\to\infty$, the integral on the right hand side is $o(\frac{1}{\sqrt{L}})$ and it follows that the contribution from the term linear in $\epsilon$ in (4.46) vanishes. Indeed the integral (4.50) can be written as
$\frac{1}{\sqrt{2L}}\int_{-L}^{L}\tilde{\epsilon}(\xi)f(xs(\xi))d\xi=\frac{1}{\sqrt{2L}}\int_{-L}^{L}\tilde{\epsilon}(\xi)\left(f(xs(\xi))-f(0)\right)d\xi+\frac{1}{\sqrt{2L}}\int_{-L}^{L}\tilde{\epsilon}(\xi)f(0)d\xi.$ (4.51)
The last integral is equal to zero in the limit $L\to\infty$ due to condition (4.45). For the other integral, assume that $f(x)$ is Lipschitz at $x=0$. Then one has, for some constant $C$, $|f(xs(\xi))-f(0)|\leq C|xs(\xi)|$, giving
$\frac{1}{\sqrt{2L}}|\int_{-L}^{L}\tilde{\epsilon}(\xi)\left(f(xs(\xi))-f(0)\right)d\xi|\leq\frac{C|x|}{\sqrt{2L}}\int_{-L}^{L}|s(\xi)|d\xi$ (4.52)
and since $s(\xi)$ is a rapidly decreasing function the left hand side in (4.52) approaches to zero in the limit $L\to\infty$ $\square$.
@@
$\dot{Q}=P.$ (4.53)
For the equation of motion for $P$ we can use Proposition (4.3), still valid if $\epsilon(\xi)$ is given as in (4.44). We get
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi.$ (4.54)
@@
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(Q\eta(\xi))\cosh(\kappa(\xi))d\xi.$ (4.55)
@@
$\mathcal{H}=\frac{P^{2}}{2}-Q+\int_{-\infty}^{\infty}\exp(Q\eta(\xi))\left(\cosh(\kappa(\xi))-1\right)d\xi.$ (4.56)
As we pointed out for the discrete case, it is possible to take a flow that is a combination of the Hamiltonian containing the even Kernel $\mathcal{H}_{e}$ (4.43), and the Hamiltonian containing the odd Kernel $\mathcal{H}_{o}$ (4.56):
$\alpha\mathcal{H}_{e}+(1-\alpha)\mathcal{H}_{o}=\frac{P^{2}}{2}-Q+\int_{-\infty}^{\infty}\left(\alpha e^{Q\eta^{o}(\xi)+\kappa^{e}(\xi)}+(1-\alpha)e^{Q\eta^{e}(\xi)+\kappa^{o}(\xi)}-1\right)d\xi.$ (4.57)
where the superscripts on the sequences remind if they are odd or even.
@@
$\dot{Q}=P,\quad\dot{P}=1-\int_{-\infty}^{+\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi,$ (4.58)
where the functions $\epsilon(x)$ and $\eta(x)$ have opposite parity with respect to $\kappa(x)$. The above system has always just one equilibrium position, say $(Q,P)=(Q_{0},0)$, if the function $\eta(x)$ has not a definite sign. The motion is bounded and the equilibrium position is a center.
Again, the function defining $\dot{P}$ in (4.58) is a decreasing function of $Q$ as it follows immediately by taking its derivative with respect to $Q$. Further, for $Q\to$-$\infty$, $\dot{P}\to+\infty$ (we assume that there will be always a set of positive measure for which $\eta(x)$ is negative), whereas for $Q\to+\infty$, $\dot{P}\to$-$\infty$ (we assume that there will be always a set of positive measure for which $\eta(x)$ is positive), and one has just one zero. The motion must be bounded for the same reasons given in Remark (3.6). Further, since (4.58) is Hamiltonian with the Hamiltonian given by
$\mathcal{H}=T+V=\frac{P^{2}}{2}-Q+\int_{-\infty}^{+\infty}\left(\exp(Q\eta(\xi)+\kappa(\xi))-1\right)dx,$ (4.59)
@@
Remark 4.6
If the function $\eta(x)$ has just one sign, then the system (4.58) has one equilibrium position if $\eta(x)\geq 0$ and has no equilibrium position (open orbits) if $\eta(x)\leq 0$.
The previous Remark follows by looking at the properties of the system (4.58) when $\eta(x)\geq 0$ or $\eta(x)\leq 0$, as did for the Remark (4.5).
Remark 4.7
The equations of motion (4.58) hold true also if the functions $\eta(x)$ and $\kappa(x)$ belong to $L^{1}$. The Propositions (4.1), (4.3) and (4.4) indeed hold also for $\eta(x)$ and $\kappa(x)$ belonging to $L^{1}$.
An example of sequences in $L^{1}$ will be given in the subsection (5.4).

### v2

@@ 4.2 The case of an even Kernel
$\epsilon(x)=\frac{1}{\sqrt{2L}}\tilde{\epsilon}(x),\quad\eta(x)=-\eta(-x)$ (91)
@@
$\dot{Q}=\int_{-\infty}^{\infty}dx\epsilon(x)\exp[P\epsilon(x)+Q\eta(x)+\kappa(x)]$ (92)
@@
$S(P,x,L)=\int_{-L}^{L}\epsilon(\xi)\exp(P\epsilon(\xi))f(xs(\xi))d\xi,$ (93)
where $f(x)$ is a continuous function of $x$ and $\epsilon(\xi)$ is given in (91). Then, for any finite values of $x$ one has
$\lim_{L\to\infty}S(P,x,L)=Pf(0).$ (94)
Proof. By taking the Taylor expansion of the exponential we get
$S(P,x,L)=\int_{-L}^{L}\epsilon(\xi)(1+P\epsilon(\xi)+\frac{P^{2}}{2}\epsilon(\xi)+\ldots)f(xs(\xi))d\xi.$ (95)
@@
$F(x,L)=\frac{1}{2L}\int_{-L}^{L}\left(f(xs(\xi))-f(0)\right)d\xi$ (97)
@@
$\lim_{L\to\infty}\frac{1}{2L}\int_{-L}^{L}\left(f(xs(\xi))-f(0)\right)d\xi=\lim_{L\to\infty}\frac{f(xs(L))+f(xs(-L))-2f(0)}{2}=0,$ (99)
where in the last equation we use the fact that $s(\xi)\in S^{\infty}$. Put in another way, in the Taylor series of the integral of $f(xs(\xi))$ (96) around $x=0$, only the first term $f(0)$ gives a finite contribution, all the others approach zero in the limit $L\to\infty$.
From the previous result it follows that all the powers of $\epsilon$ greater than two in the expansion of $\epsilon\exp(P\epsilon)$ in (93) don’t give a contribution in the limit $L\to\infty$.
Let us look at the term linear in $\epsilon$ in (93). One has
$\int_{-L}^{L}\epsilon(\xi)f(xs(\xi))d\xi=\frac{1}{\sqrt{2L}}\int_{0}^{L}\tilde{\epsilon}\left(f(xs(\xi))-f(xs(-\xi))\right)d\xi$ (100)
@@
$\frac{1}{\sqrt{2L}}\int_{0}^{L}\tilde{\epsilon}\Big(f(xs(\xi))-f(0)\Big)d\xi-\int_{0}^{L}\tilde{\epsilon}\Big(f(xs(-\xi))-f(0)\Big)d\xi$ (101)
@@
$\begin{split}&|\int_{-L}^{L}\epsilon(\xi)f(xs(\xi))d\xi|\leq\frac{C|x|}{\sqrt{2L}}\int_{0}^{L}\left(|s(\xi)-s(-\xi)|\right)d\xi\end{split}$ (102)
@@
Corollary 4.2.
The equation of motion (92) reduces, in the limit $L\to\infty$, to
$\dot{Q}=P.$ (103)
@@
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(P\epsilon+Q\eta+\kappa)d\xi.$ (104)
@@
$T(P,x,L)=\int_{-L}^{L}\eta(\xi)\exp(P\epsilon(\xi))f(xs(\xi))d\xi,$ (105)
where $f(x)$ is a continuous function of $x$ and $\epsilon(\xi)$ is given in (91). Then, for any finite values of $x$ one has
$\lim_{L\to\infty}T(P,x,L)=\int_{-\infty}^{\infty}\eta(\xi)f(xs(\xi))d\xi$ (106)
Proof. For $\xi>\xi_{0}$, since $s(\xi)\to 0$ for $\xi\to\infty$, by continuity one has $|f(xs(\xi))-f(0)|<\delta(x)$ or $|f(xs(\xi))|<|f(0)|+\delta(x)$ So one has
$\begin{split}&|\int_{-\infty}^{\infty}\eta(\xi)f(xs(\xi))d\xi|=|\int_{|t|<t_{0}}\eta(\xi)f(xs(\xi))d\xi+\int_{|t|\geq t_{0}}\eta(\xi)f(xs(\xi))d\xi|\leq\\ &\leq\int_{|t|<t_{0}}|\eta(\xi)f(xs(\xi))|d\xi+(|f(0)|+\delta(x))\int_{|t|\geq t_{0}}|\eta(\xi)|d\xi,\end{split}$ (107)
i.e. the integral $\int_{-\infty}^{\infty}\eta(\xi)f(xs(\xi))d\xi$ is finite since $\eta(\xi)$ is in $S^{\infty}$. It follows that if one considers the Taylor expansion of $\exp(P\epsilon)$ in the integral (105), only the term proportional to $\epsilon^{0}$ survives in the limit $L\to\infty$ and we get (106). $\square$
@@
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi.$ (108)
@@
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(\kappa(\xi))\sinh(Q\eta(\xi))d\xi.$ (109)
@@
$\mathcal{H}=\frac{P^{2}}{2}-Q+\int_{-\infty}^{\infty}\exp(\kappa(\xi))\left(\cosh(Q\eta(\xi))-1\right)d\xi$ (110)
@@ 4.3 The case of an odd Kernel
$\epsilon(x)=\frac{1}{\sqrt{2L}}\tilde{\epsilon}(x),\quad\eta(x)=\eta(-x)$ (111)
@@
$\lim_{L\to\infty}\int_{-L}^{L}\epsilon(\xi)d\xi=o(\sqrt{L}).$ (112)
@@
$M(P,x,L)=\int_{-L}^{L}\epsilon(\xi)\exp(P\epsilon(\xi))f(xs(\xi))d\xi,$ (113)
where $f(x)$ is a continuous function of $x$ and $\epsilon(\xi)$ is given in (111) and satisfies (112). Then, for any finite values of $x$ one has
$\lim_{L\to\infty}M(P,x,L)=Pf(0).$ (114)
Proof. By taking the Taylor series of the exponential in (113) one has
$M(P,x,L)=\int_{-L}^{L}\epsilon(\xi)\left(1+P\epsilon(\xi)+\frac{P^{2}}{2}\epsilon(\xi)^{2}+\ldots\right)f(xs(\xi))d\xi.$ (115)
@@
$P\int_{-L}^{L}\epsilon(\xi)^{2}f(xs(\xi))d\xi=\frac{1}{2L}\int_{-L}^{L}f(xs(\xi))d\xi.$ (116)
Exactly as for the equation (96), the previous expressions equals $Pf(0)$. It remains only the term linear in $\epsilon$ to look at, since all the powers of $\epsilon$ greater than $2$ in the Taylor expansion (115) do not contribute in the limit $L\to\infty$. The linear term in $\epsilon$ is
$\int_{-L}^{L}\epsilon(\xi)f(xs(\xi))d\xi=\frac{1}{\sqrt{2L}}\int_{-L}^{L}\tilde{\epsilon}(\xi)f(xs(\xi))d\xi.$ (117)
Due to the condition (112), since $f(xs(\xi))\to f(0)$ as $\xi\to\infty$, the integral on the right hand side is $o(\frac{1}{\sqrt{L}})$ and it follows that the contribution from the term linear in $\epsilon$ in (113) vanishes. Indeed the integral (117) can be written as
$\frac{1}{\sqrt{2L}}\int_{-L}^{L}\tilde{\epsilon}(\xi)f(xs(\xi))d\xi=\frac{1}{\sqrt{2L}}\int_{-L}^{L}\tilde{\epsilon}(\xi)\left(f(xs(\xi))-f(0)\right)d\xi+\frac{1}{\sqrt{2L}}\int_{-L}^{L}\tilde{\epsilon}(\xi)f(0)d\xi.$ (118)
The last integral is equal to zero in the limit $L\to\infty$ due to condition (112). For the other integral, assume that $f(x)$ is Lipschitz at $x=0$. Then one has, for some constant $C$, $|f(xs(\xi))-f(0)|\leq C|xs(\xi)|$, giving
$\frac{1}{\sqrt{2L}}|\int_{-L}^{L}\tilde{\epsilon}(\xi)\left(f(xs(\xi))-f(0)\right)d\xi|\leq\frac{C|x|}{\sqrt{2L}}\int_{-L}^{L}|s(\xi)|d\xi$ (119)
and since $s(\xi)$ is a rapidly decreasing function the left hand side in (119) approaches to zero in the limit $L\to\infty$ $\square$.
@@
$\dot{Q}=P.$ (120)
For the equation of motion for $P$ we can use Proposition (4.3), still valid if $\epsilon(\xi)$ is given as in (111). We get
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi.$ (121)
@@
$\dot{P}=1-\int_{-\infty}^{\infty}\eta(\xi)\exp(Q\eta(\xi))\cosh(\kappa(\xi))d\xi.$ (122)
@@
$\mathcal{H}=\frac{P^{2}}{2}-Q+\int_{-\infty}^{\infty}\exp(Q\eta(\xi))\left(\cosh(\kappa(\xi))-1\right)d\xi.$ (123)
@@
$\dot{Q}=P,\quad\dot{P}=1-\int_{-\infty}^{+\infty}\eta(\xi)\exp(Q\eta(\xi)+\kappa(\xi))d\xi,$ (124)
where the functions $\epsilon(x)$ and $\eta(x)$ have opposite parity with respect to $\kappa(x)$. The above system has always just one equilibrium position, say $(Q,P)=(Q_{0},0)$, if the function $\eta(x)$ has not a definite sign. In this case the motion is bounded and the equilibrium position is a center.
Again, the function defining $\dot{P}$ in (124) is a decreasing function of $Q$ as it follows immediately by taking its derivative with respect to $Q$. Further, for $Q\to$-$\infty$, $\dot{P}\to+\infty$ (we assume that there will be always a set of positive measure for which $\eta(x)$ is negative), whereas for $Q\to+\infty$, $\dot{P}\to$-$\infty$ (we assume that there will be always a set of positive measure for which $\eta(x)$ is positive), and one has just one zero. The motion must be bounded for the same reasons given in Remark (3.6). Further, since (124) is Hamiltonian with the Hamiltonian given by
$\mathcal{H}=T+V=\frac{P^{2}}{2}-Q+\int_{-\infty}^{+\infty}\left(\exp(Q\eta(\xi)+\kappa(\xi))-1\right)dx,$ (125)
@@
Remark 4.6.
If the function $\eta(x)$ has just one sign, then the system (124) has one equilibrium position if $\eta(x)\geq 0$ and has no equilibrium position (open orbits) if $\eta(x)\leq 0$.
The previous Remark follows by looking at the properties of the system (124) when $\eta(x)\geq 0$ or $\eta(x)\leq 0$, as did for the Remark (4.5).
Remark 4.7.
The equations of motion (124) hold true also if the functions $\eta(x)$ and $\kappa(x)$ belong to $L^{1}$. The Propositions (4.1), (4.3) and (4.4) indeed hold also for $\eta(x)$ and $\kappa(x)$ belonging to $L^{1}$.
An example of sequences in $L^{1}$ will be given in the subsection (5.4). We end this Section giving a Remark, like in the discrete case, about the possibility to get a mixed odd/even Hamiltonian.
Remark 4.8.
It is possible to take a flow that is a combination of the Hamiltonian containing the even Kernel $\mathcal{H}_{e}$ (110), and the Hamiltonian containing the odd Kernel $\mathcal{H}_{o}$ (123):
$\alpha\mathcal{H}_{e}+(1-\alpha)\mathcal{H}_{o}=\frac{P^{2}}{2}-Q+\int_{-\infty}^{\infty}\left(\alpha e^{Q\eta^{o}(\xi)+\kappa^{e}(\xi)}+(1-\alpha)e^{Q\eta^{e}(\xi)+\kappa^{o}(\xi)}-1\right)d\xi.$ (126)
where the superscripts on the sequences remind if they are odd or even. If the functions $\eta^{o}(\xi)$ and $\eta^{e}(\xi)$ have not a definite sign, the potential
$V=-Q+\int_{-\infty}^{\infty}\left(\alpha e^{Q\eta^{o}(\xi)+\kappa^{e}(\xi)}+(1-\alpha)e^{Q\eta^{e}(\xi)+\kappa^{o}(\xi)}-1\right)d\xi$ (127)
has just one minima for $\alpha\in[0,1]$ and is an infinite potential well. so that the corresponding motion is bounded and periodic. The proof being the same as in Remark (4.5) and is omitted.